OpenStudy (baldymcgee6):

Can someone explain this? sqrt(x+3) + x +3 = sqrt(x+1)[1+sqrt(x+3)]

OpenStudy (baldymcgee6):

\[\sqrt{x+3} + x + 3 = \sqrt{x+1} (1+\sqrt{x+3})\]

OpenStudy (anonymous):

This is an eqn - open the brackets and start raising in 2-nd degree

OpenStudy (baldymcgee6):

I'm just trying to figure out how they factored this

OpenStudy (anonymous):

\[\sqrt{(.....) * (...)} =\sqrt{(.....)}\sqrt{(...)}\]

OpenStudy (anonymous):

Only you made a small typo : it should be:\[\sqrt{x+3} + x + 3 = \sqrt{x+3}(1 + \sqrt{x+3})\]

OpenStudy (baldymcgee6):

ahhh, I did not make the typo, that is straight out of a textbook.

OpenStudy (baldymcgee6):

even so, how do you pull out an sqrt(x+3) from x?

OpenStudy (anonymous):

See the rule of root multiplication that I posted for u above - the content "merges" and multiplies

OpenStudy (anonymous):

start by dividing everything by \[\sqrt{x+3}\]

OpenStudy (baldymcgee6):

@Mikael, but we are not multiplying anything to begin with...?

OpenStudy (anonymous):

\[x+3=\sqrt{(x+3)*(x+3)}=\sqrt{(x+3)}*\sqrt{(x+3)}\]

OpenStudy (anonymous):

You do multiply - ON THE LEFT HAND SIDE !

OpenStudy (anonymous):

Sorry RIGHT HAND SIDE

OpenStudy (anonymous):

OPEN the parentheses - u have to multiply

OpenStudy (baldymcgee6):

@juantweaver how does x+3=sqrt((x+3)∗(x+3))

OpenStudy (anonymous):

another example\[2=\sqrt{2*2}=\sqrt{2}*\sqrt{2}\]

OpenStudy (anonymous):

Come ON\[\sqrt{A}*\sqrt{A} = \sqrt{A^2} = A\]

OpenStudy (baldymcgee6):

oohh k thanks

OpenStudy (anonymous):

:}

OpenStudy (anonymous):

close the question pl

OpenStudy (anonymous):

koala vs. gator :P

OpenStudy (anonymous):

No we live indifferent Eco-niches and therefore are smiling at each other :]]]

OpenStudy (anonymous):

I am on top (of the forest) and gator is ... hmm, well you know...

OpenStudy (anonymous):

Please close the question @baldymcgee6

OpenStudy (anonymous):

Please close the question @baldymcgee6

OpenStudy (baldymcgee6):

sorry, battery died.

OpenStudy (anonymous):

Thanks for "fan" meet u again but also u have to read beforehand u ask

OpenStudy (baldymcgee6):

Read what before I ask?